the length of the rectangle is 9 cm more than its width.Its area is 112 square.find the dimension of the rectangle
Answer by Cromlix(4381)

Width = x
Length = x + 9
Area = Length * Width
112 = (x + 9) * x
112 = x^2 + 9x
x^2 + 9x - 112 = 0
(x - 7)(x + 16) = 0
x + 16 = 0
x = -16 (no answer as x < 0)
x - 7 = 0
x = 7
Width = 7cm
Length = 16cm
Hope this helps.
:-)